OpenStudy (anonymous):

For a triangle with sides a, b and c, what is the range of the length of side c? a = 620, b = 1050, c = < c <

OpenStudy (anonymous):

hey can you comeback to mines plz

OpenStudy (anonymous):

lol umm which one

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

lol hold on

OpenStudy (anonymous):

ok this one is hard for me sorry maybe fiddle can help u

OpenStudy (kira_yamato):

Use the cosine rule... The lower limit is when the angle is 0 radians, and the upper limit is when the angle is π radians

OpenStudy (kira_yamato):

\[c^2 = a^2 + b^2 -2ab \cos \theta\]